# AI red-teaming is a sociotechnical challenge: on values, labor, and harms

### v2

@@ 4. Well-being of Red-Teamers
Beyond understanding who is doing the AI red-teaming and what is being evaluated, we also need to pay attention to the human cost of doing such work. Scholars and practitioners involved in red-teaming call it “rather unsavory work” (Bai et al., 2022) and “mentally taxing” (Microsoft, 2023). Like content moderation work, the adversarial probing of red-teaming requires workers to imagine worst-case scenarios and expose themselves to potentially troubling outputs. Red-teamers may be required to (1) assume the persona of a potential adversary (e.g., online harasser, sex trafficker, racist, terrorist), (2) invent a plan that this adversary may use to compromise the system, and (3) evaluate the output for potential harms. They may also assume the persona of a benign user with specific intents or contexts (e.g., a user with a history of eating disorders looking for dieting advice) and try to reveal system vulnerabilities that might be harmful. A red-teamer may have to first research harmful groups to learn their behaviors and bring that knowledge into red-teaming the models for hate speech or deep fakes. They may immerse themselves in child online safety concerns to then evaluate the model’s capabilities in aiding child exploitation.
@@
To date, there is little empirical research about the psychological impact of AI red-teaming. For example, one recent study of RAI content workers shows that red-teamers, content moderators, and data labelers face similar psychological challenges when working with potentially harmful content (Zhang et al., 2024a). Given the extensive research on content moderation and the well-documented occupational health concerns experienced by professions that contend with trauma exposure (Steiger et al., 2021), red-teamers stand to benefit from this history.
@@
Moreover, the success of a red-team operation depends on uncovering and reviewing increasingly harmful content. Much like content moderators, emergency responders, journalists, or police investigators dealing with distressing events, AI red-teamers may experience repeated exposure to disturbing and traumatic content that can lead to negative psychological symptoms. For example, content moderators removing harmful and offensive material from platforms have reported mental health challenges that extend long after the work is complete, leading to documented cases of post-traumatic stress disorder (PTSD) and secondary traumatic stress (STS) (Ruckenstein and Turunen, 2019; Arsht and Etcovitch, 2018). Prolonged exposure may contribute to long-lasting mental health symptoms, alterations to their personal belief systems, and increased risks of physical health issues and substance abuse (Steiger et al., 2021). In fact, repeated work-related exposure to traumatic content is among the diagnostic criteria used for PTSD in the DSM5 (Diagnostic and Statistical Manual of Mental Disorders) and has subsequently been used to support the claims in a series of recent lawsuits brought on by content moderators against their employers (Pinchevski, 2023). [^14] See https://www.nytimes.com/2018/09/25/technology/facebook-moderator-job-ptsd-lawsuit.html Although many major platforms were slow to deal with this problem, most now offer mental health support and take measures to limit moderators’ exposure to the most reprehensible content. These parallels underscore the importance of initiating research to protect red-teamers from the psychological hazards inherent in their work.
@@
At the same time, AI red-teaming introduces distinct psychological challenges. A successful AI red-teamer must exhibit an antagonistic imagination to be effective. Or as one red-teamer put it: “If there were a red-team motto, it would be: The more sinister your imagination, the better your work.” [^15] https://www.bostonglobe.com/2024/01/11/opinion/ai-testing-red-team-human-toll/ Red-teaming involves deliberately engaging in transgressive, uncomfortable, unethical, immoral, or harmful activities, including immersing themselves in scenarios that go against their morals or belief systems–to think like a harasser, or feel like a target of discrimination. As documented in (Zhang et al., 2024a), such practice can lead to “moral injury” (Shay, 2014), a form of psychological distress that stems from actions, or the lack thereof, that violate one’s moral or ethical code. Those who cannot safely detach their personal identity from their transgressions may experience negative self-perception and guilt. Regularly breaking the rules for the greater good can introduce a potential for a “loss of self,” sometimes seen in the undercover police profession (Joh, 2009).
@@
The potential negative impacts on red-teamers’ wellbeing have been acknowledged by those that organize such work. For example, organizers of the DEFCON Generative Red Team event anticipated that models generating unexpected harmful outputs might be triggering to participants (Cattell, 2023). Anthropic’s early red-teaming efforts involved consultations with Trust & Safety professionals to design safety measures for their crowdworkers (Ganguli et al., 2022). Strategies for preserving the wellbeing of red-teamers could include providing warnings about sensitive content, allowing opt-outs, encouraging breaks, monitoring mood, or allowing them to choose topics within their own risk tolerance (Ganguli et al., 2022; noa, 2023; Microsoft, 2023). But for volunteer or crowdsourced red-teamers, such strategies are limited to what the organizers are willing to provide.
@@
For professional red-teamers, companies sometimes offer employee assistance programs (EAPs) with mental health resources. However, organizational factors often impact how these resources are actually used. For some workers, accessing therapists may be a luxury they cannot afford, because of a psychologically unsafe work environment, unrelenting performance metrics, or job insecurity. Non-disclosure agreements (NDAs) have historically prevented workers from speaking up about working conditions (Steiger et al., 2021). Monitoring red-teamers’ wellbeing gets entangled with more unsavory forms of workplace surveillance (Ajunwa et al., 2017), and increases liability risks for employers if findings are severely negative. And implementing consistent wellbeing strategies can become infeasible for red-teamers working across organizational and national boundaries. Barriers to accessing mental health care, for both paid and unpaid red-teamers, include not only a shortage of providers but also stigma and lack of awareness (Haugen et al., 2017). So when organizers claim red-teaming events were safe because no one used the on-call therapists (Cattell, 2023), they may be mistakenly assuming that no usage means no need.
@@
In an attempt to minimize human exposure to the “unsavory work” of red-teaming, some researchers have advocated for automated red-teaming. Whether or not this is possible–and some leading AI safety researchers have recently argued that the ”human element” will always be necessary for red-teaming (Bullwinkel et al., 2025)–automation may inadvertently make the human work even more invisible (Gray and Suri, 2019), diverting resources away from well-being measures needed by the red-teamers who remain. Therefore, it is important to critically examine the role of automation in red-teaming—not only its immediate impact on tasks, but also its long-term effects on the overall ecosystem.
We must acknowledge the sobering reality that accompanies the commoditization of AI harm reduction. As long as generative AI remains integral to our lives, the work of AI red-teaming and its psychological implications will persist. To support this workforce, it is crucial to rigorously study and validate the effectiveness of innovative well-being strategies across various contexts, with close examination of the surrounding organizational and social structures.
@@ 5. Conclusion
What would more substantive, empirical, and cross-disciplinary research on red-teaming provide? Today’s siloed, firewalled, market-reactive approach to red-teaming has potential drawbacks for AI’s consumers, red-teamers, and AI companies. Each company is rapidly developing its own version of red-teaming, with definitions and workloads varying based on the company’s priorities, ‘brand,’ and particular focus. Almost every company working on generative AI today has a red team workforce of some kind. While there is energy being put toward addressing biases and other “embedded harms,” plenty of red-teaming efforts are more concerned with ungrounded, existential risks than with current, tangible concerns. It is even unclear how many issues identified through red-teaming efforts have been mitigated.
An organized, empirical research agenda that examines red-teaming as a sociotechnical undertaking could make both AI companies and the public more aware of and intentional about the practices and consequences of red-teaming work. Regardless of its current effectiveness or its future improvements, red-teaming has underlying logics and structural conditions that need examination. Lessons from content moderation show that finding and removing ”bad elements” demands the deliberations and value judgments of teams of people. The specific conditions under which people do this hard work, at an unprecedented global scale and across myriad institutional settings, matters for both red team workers’ occupational health and for the integrity of our technical and informational ecosystems.
We must study how red-teaming judgments are made and by whom if we hope to improve its outcomes as a sociotechnical system. Its compartmentalization and operational opacity can both alienate workers and keep the public from understanding fully what AI systems offer. Empirical study can challenge these arrangements, identify the barriers and incentives at play, and perhaps point the public and AI companies to more sustainable alternatives.
And like most data work, the notion that this labor is only temporary, that it will soon be automated away, is wrong, and (deliberately) distracting from these sociological concerns. It is difficult to believe that we can ever fully automate such peculiarly human judgments, about contentious and shifting topics, under pressure from regulators and the public, whose ethical frameworks can themselves shift. But even if it could be automated in the future, real people are doing this work right now, and with real consequences. It makes little difference whether we discard this phantasmic notion of full automation entirely, or just concede that data labor will be with us for the foreseeable future. Either way, research today can help identify how to structure this work in ways that are attentive to the well-being and the labor rights of the people doing the work right now.

### v3

@@ 4. Well-being of Red-Teamers
Beyond understanding who is doing the AI red-teaming and what is being evaluated, we also need to pay attention to the human cost of doing such work. Scholars and practitioners involved in red-teaming call it “rather unsavory work” (Bai et al., 2022) and “mentally taxing” (Microsoft, 2023). Like content moderation work, the adversarial probing of red-teaming requires workers to imagine worst-case scenarios and expose themselves to potentially troubling outputs. Red-teamers are often required to assume the persona of a potential adversary (for example: online harasser, sex trafficker, racist, terrorist), invent a plan that this adversary may use to compromise the system, and evaluate the output for potential harms. They may also assume the persona of a benign user with specific intents or contexts (for example, a user with a history of eating disorders looking for dieting advice) and try to reveal system vulnerabilities that might be harmful. A red-teamer may first research harmful groups to learn their behaviors and bring that knowledge into red-teaming the models for hate speech or deep fakes. They may immerse themselves in child online safety concerns to then evaluate the model’s capabilities in aiding child exploitation.
@@
To date, there is little empirical research about the psychological impact of AI red-teaming. But one recent study of RAI content workers shows that red-teamers, content moderators, and data labelers face similar psychological challenges when working with potentially harmful content (Zhang et al., 2024a). Given the extensive research on content moderation and the well-documented occupational health concerns experienced by professions that contend with trauma exposure (Steiger et al., 2021), red-teamers stand to benefit from attention to this history.
@@
Moreover, the success of a red-team operation depends on uncovering and reviewing increasingly harmful content. Much like content moderators, emergency responders, journalists, or police investigating distressing events, AI red-teamers may be exposed repeatedly to disturbing and traumatic content that can lead to negative psychological symptoms. For example, content moderators removing harmful and offensive material from platforms have reported mental health challenges extending long after the work is complete, leading to documented cases of post-traumatic stress disorder (PTSD) and secondary traumatic stress (STS) (Ruckenstein and Turunen, 2019; Arsht and Etcovitch, 2018). Prolonged exposure may contribute to long-lasting mental health symptoms, alterations to their personal belief systems, and increased risks of physical health issues and substance abuse (Steiger et al., 2021). In fact, repeated work-related exposure to traumatic content is among the diagnostic criteria used for PTSD in the DSM5 (Diagnostic and Statistical Manual of Mental Disorders) and has subsequently been used to support a series of recent lawsuits brought by content moderators against their employers (Pinchevski, 2023). [^14] See https://www.nytimes.com/2018/09/25/technology/facebook-moderator-job-ptsd-lawsuit.html Although many major platforms were slow to deal with this problem, most now offer mental health support and take measures to limit moderators’ exposure to the most reprehensible content. These parallels underscore the importance of initiating research to protect red-teamers from the psychological hazards inherent in their work.
@@
At the same time, AI red-teaming introduces distinct psychological challenges. A successful AI red-teamer must exhibit an antagonistic imagination to be effective. Or as one red-teamer put it: “If there were a red-team motto, it would be: The more sinister your imagination, the better your work.” [^15] https://www.bostonglobe.com/2024/01/11/opinion/ai-testing-red-team-human-toll/ Red-teaming involves deliberately engaging in transgressive, uncomfortable, unethical, immoral, or harmful activities, including immersing themselves in scenarios that go against their morals or belief systems–to think like a harasser, or feel like a target of discrimination. Such practice can lead to “moral injury” (Shay, 2014), a form of psychological distress that stems from actions, or the lack thereof, that violate one’s moral or ethical code (Zhang et al., 2024a). Those who cannot safely detach their personal identity from their transgressions may experience negative self-perception and guilt. Regularly breaking the rules for the greater good can lead to a “loss of self,” sometimes seen in the undercover police profession (Joh, 2009).
@@
The potential negative impacts on red-teamers’ wellbeing have been acknowledged by some of those who organize such work. For example, organizers of the DEFCON Generative Red Team event anticipated that models generating unexpected harmful outputs might be triggering to participants (Cattell, 2023). Anthropic’s early red-teaming efforts involved consultations with Trust & Safety professionals to design safety measures for their crowdworkers (Ganguli et al., 2022). Strategies for preserving the wellbeing of red-teamers could include providing warnings about sensitive content, allowing opt-outs, encouraging breaks, monitoring mood, or allowing them to choose topics within their own risk tolerance (Ganguli et al., 2022; OpenAI, 2023; Microsoft, 2023). But for volunteer or crowdsourced red-teamers, such strategies are limited to what the organizers are willing to provide.
@@
For professional red-teamers, companies sometimes offer employee assistance programs (EAPs) with mental health resources. However, organizational factors often impact how these resources are actually used. For some, accessing a therapist may be a luxury they cannot afford, because of a psychologically unsafe work environment, unrelenting performance metrics, or job insecurity. Non-disclosure agreements (NDAs) have historically prevented workers from speaking up about working conditions (Steiger et al., 2021). Monitoring red-teamers’ wellbeing gets entangled with more unsavory forms of workplace surveillance (Ajunwa et al., 2017), and increases liability risks for employers if the findings are severely negative. And implementing consistent wellbeing strategies can become infeasible for red-teamers working across organizational and national boundaries. So when organizers claim red-teaming events were safe because no one used the on-call therapists (Cattell, 2023), they may be mistakenly assuming that no usage means no need.
@@
In an attempt to minimize human exposure to the “unsavory work” of red-teaming, some researchers have advocated for automated red-teaming. Whether or not this is possible–some leading AI safety researchers have recently argued that the ”human element” will always be necessary (Bullwinkel et al., 2025)–automation may inadvertently make the human work even more invisible (Gray and Suri, 2019), diverting resources away from well-being measures needed by those red-teamers who remain. Therefore, it is important to critically examine the role of automation in red-teaming—not only its immediate impact on tasks, but also its long-term effects on the overall ecosystem.
We must acknowledge the sobering reality that comes with the commoditization of AI harm reduction. As long as generative AI remains integral to our lives, the work of AI red-teaming and its psychological implications will persist. To support this workforce, it is crucial to rigorously study and validate the effectiveness of innovative well-being strategies across various contexts, with close examination of the surrounding organizational and social structures.
@@ 5. Conclusion
What would more substantive, empirical, and cross-disciplinary research on red-teaming provide? Today’s siloed, firewalled, market-reactive approach to red-teaming has potential drawbacks for AI consumers, red-teamers, and companies. Each company is rapidly developing its own version of red-teaming, with definitions and workloads varying based on the company’s priorities, ‘brand,’ and particular focus. Almost every company working on generative AI today has a red team workforce of some kind. While there is energy being put toward addressing biases and other “embedded harms,” plenty of red-teaming efforts are more concerned with ungrounded, existential risks than with current, tangible concerns. And it is unclear how many issues identified through red-teaming efforts have been mitigated.
We must study how red-teaming judgments are made and by whom if we hope to improve its outcomes as part of a sociotechnical system. Its compartmentalization and operational opacity can both alienate workers and keep the public from understanding fully what AI systems can offer. Empirical study can challenge these arrangements, identify the barriers and incentives at play, and perhaps point the public and AI companies toward more sustainable alternatives.
An organized, crossdisciplinary, empirical research agenda that examines red-teaming as a sociotechnical undertaking could make both AI companies and the public more aware of and intentional about red-teaming work. Regardless of its current effectiveness or its future improvements, red-teaming has underlying logics and structural conditions that need examination. Lessons from content moderation show that finding and removing ”bad elements” demands the deliberations and value judgments of teams of people. The specific conditions under which people do this hard work, at an unprecedented global scale and across myriad institutional settings, matters for both red team workers’ occupational health and for the integrity of our technical and informational ecosystems.
And like most data work, the notion that this labor is only temporary, that it will soon be automated away, is wrong, and (deliberately) distracting from these sociological concerns. It is difficult to believe that we can ever fully automate such peculiarly human judgments, about contentious and shifting topics, under pressure from regulators and the public, whose ethical frameworks can themselves shift. But even if it could be automated in the future, real people are doing this work right now, and with real consequences. It makes little difference whether we discard this phantasmic notion of full automation entirely, or just concede that data labor will be with us for the foreseeable future. Either way, research today can help structure this work in ways that are better attentive to the well-being and labor rights of the people doing the work right now.